### v1

# Improving Cramér–Rao Bound And Its Variants: An Extrinsic Geometry Perspective

Abstract
This work presents a geometric refinement of the classical Cramér–Rao bound (CRB) by incorporating curvature-aware corrections based on the second fundamental form associated with the statistical model manifold. That is, our formulation shows that relying on the extrinsic geometry of the square root embedding of the manifold in the ambient Hilbert space comprising square integrable functions with respect to a fixed base measure offers a rigorous (and intuitive) way to improve upon the CRB and some of its variants, such as the Bhattacharyya-type bounds, that use higher-order derivatives of the log-likelihood. Precisely, the improved bounds in the latter case make explicit use of the elegant framework offered by employing the Faà di Bruno formula and exponential Bell polynomials in expressing the jets associated with the square root embedding in terms of the raw scores. The interplay between the geometry of the statistical embedding and the behavior of the estimator variance is quantitatively analyzed in concrete examples, showing that our corrections can meaningfully tighten the lower bound, suggesting further exploration into connections with estimator efficiency in more general situations.
keywords:
Bhattacharyya Bound, Cramér-Rao Bound, Curvature, Second Fundamental Form, Statistical Manifold, Square Root Embedding.
1 Introduction
The Cramér-Rao Bound (CRB) is a foundational result in parametric estimation theory, specifying a lower limit on the variance of any (locally) unbiased estimator under appropriate regularity conditions, and underpins the concept of estimator efficiency [1, 2, 3]. While widely used, the CRB may be loose in several practical settings – such as nonlinear models, low signal-to-noise regimes, or finite sample sizes – leading to extensive literature on improved lower bounds. Notable classical refinements include the Chapman-Robbins bound [4] and the Ziv-Zakai bound [5], which generalize the CRB beyond the usual regularity assumptions, such as differentiability under the integral sign, or invoke optimal detection problems. Bayesian and hybrid variants of the CRB have been developed to incorporate prior uncertainty that avoid costly matrix inversions [6]. Computational saving was also achieved with constrained CRBs and singular Fisher information matrices in [7].
A line of work that is of relevance to us includes the classical Bhattacharyya inequality ([8]; see also [9, 10]) that generalizes the CRB by incorporating higher-order derivatives of the log-likelihood and is of importance when an estimator that is efficient in the sense of the CRB is not available. While clarifying the algebraic structure of the Bhattacharyya bound, a hierarchy of improvements – an increasing sequence of variance bounds that converge to the estimator variance – was shown by Pillai and Sinha [11] based on projections of the estimator error onto spans of higher-order derivatives of the log-likelihood. This elegant formulation of the Bhattacharyya bound, however, lacks a geometric interpretation despite its Hilbert space foundation. In short, these constructions remain intrinsically score-based and extrinsic geometry of the embedded statistical manifold, which results in nontrivial correction if the estimator error does not lie in the span of the higher-order derivatives considered, is not accounted for; this is exactly what we point out in our research.
Information geometry offers an alternative framework in which statistical models are viewed as Riemannian manifolds endowed with the Fisher-Rao metric. This viewpoint allows estimator performance to be studied using geometric concepts such as curvature, geodesics, and connections [12]. Nevertheless, most information geometry-based works tend to focus on intrinsic properties and divergence measures, as briefly discussed below.
Several works have extended the CRB to intrinsic Riemannian settings. In an important contribution, Smith [13] derived intrinsic versions of the CRB on manifolds utilizing an arbitrary affine connection with arbitrary geodesics for both biased and unbiased estimators while Barrau and Bonnabel [14], in a follow-up, elucidated the extent to which the intrinsic CRB obtained by Smith is independent of a specific choice of coordinates. Further, the authors also addressed a question revolving around the dependence of the intrinsic bound on the underlying parameter. Extending these results to the Bayesian setting, Bouchard et al. [15] studied the problem of covariance matrix estimation when the data is sampled from a Gaussian distribution whose covariance matrix is drawn from an inverse Wishart distribution. Performance bounds for this problem were obtained, for both the mean squared error (Euclidean metric) and the natural Riemannian distance for Hermitian positive definite matrices (affine invariant metric), using the derived intrinsic Bayesian CRB. Interestingly, it was shown that the maximum a posteriori and the minimum mean squared error estimators are asymptotically efficient in the intrinsic case, and not so in the Euclidean situation.
More recently, building upon the work of Li and Zhao [16] who derived the Wasserstein–Cramér-Rao inequality – a lower bound on the Wasserstein variance of an estimator, Nishimori and Matsuda [17] investigated conditions for an estimator to attain the lower bound asymptotically. In an interesting application to chemical reaction networks, Loutchko et al. [18] developed, among various other results, a multivariate CRB through the comparison of two Riemannian metric tensors.
Another important thread involves divergence-based generalizations of the CRB. Mishra and Kumar [19] proposed a general Riemannian metric based on relative $\alpha$-entropy to obtain a generalized Bayesian Cramér-Rao inequality, thereby establishing a lower bound for the variance of an unbiased estimator for the $\alpha$-escort distribution starting from an unbiased estimator for the underlying distribution. Subsequently, Kumar and Mishra [20] provided a unifying differential geometric framework for deriving CRB-like inequalities from generalized Csiszár divergences. They applied Eguchi’s theory to derive the Fisher information metric and dual affine connections arising from these generalized divergence functions, which enabled them to prove a more widely applicable version of the Cramér-Rao inequality. In [21], these authors again relied on Eguchi’s theory of obtaining dualistic geometric structures from a divergence function and then applied Amari-Nagoaka’s theory [12] to obtain a Cramér-Rao type inequality. These approaches serve to find unbiased and efficient estimators, but typically remain in an intrinsic, divergence-induced geometry.
However, we note that a conceptual gap remains. The algebraic projection-based CRBs, such as those of Bhattacharyya [8]/Pillai and Sinha [11], do not exploit the geometric curvature of the statistical manifold. While it may be argued that the second order Bhattacharyya-type bounds also incorporate second derivatives of the density and hence (in some sense) capture second-order information, these do not directly/correctly capture the inherent geometry via the curvature information. Conversely, information geometry-based CRBs focus on intrinsic structures and rarely intersect with projection-theoretic Hilbert space methods. In particular, there is little work incorporating extrinsic geometric effects – such as the second fundamental form – into variance bounds.
In this work, we bridge this gap by developing a Hilbert space projection-based and Riemannian geometric refinement of classical and higher-order CRBs. Suppose the family of probability distributions $\{P_{\theta}\}$, parameterized by the scalar $\theta\in\Theta$, is absolutely continuous with respect to a fixed base measure $\mu$ with a strictly positive density $f(\cdot;\theta)$. Define the square root embedding $s_{\theta}:\Theta\mapsto L^{2}(\mu)$ by $s_{\theta}:=\sqrt{f(\cdot;\theta)}$, where $L^{2}(\mu)$ is the space of square integrable functions with respect to $\mu$, and consider the jets
$\eta_{k}(\theta):=\partial_{\theta}^{k}s_{\theta},\qquad k\geq 1.$
Starting from a projection-theoretic decomposition of estimator error in the ambient space $L^{2}(\mu)$, we derive a residual component orthogonal to the span of $\eta_{1}$. We then show that the contribution of this residual can be explicitly quantified using the second fundamental form of the statistical manifold embedded in $L^{2}(\mu)$, yielding a curvature-aware correction to the CRB. A variance bound that explicitly betters the Bhattacharyya/Pillai–Sinha-type results is then derived by further analyzing the structure of the $\eta_{k}$ in relation to the true (raw) score and its derivatives $\partial_{\theta}^{k}\log f(\cdot;\theta)$. Importantly, note that we refer to results that use $\partial_{\theta}^{k}\log f(\cdot;\theta)$ for $k\geq 1$ as Bhattacharyya-type (or Pillai–Sinha-type) bounds throughout the paper since the classical Bhattacharyya bound used $\frac{\partial^{k}_{\theta}f}{f}$ instead.
Our contributions are as follows:
  • We derive extrinsic geometric refinement terms using the second fundamental form, providing new insight into how curvature controls estimator error beyond linear approximations, and establish improvement over the classical CRB – see Theorem 2.
  • We interpret the higher-order variance lower bound sequence obtained by Bhattacharyya [8]/Pillai and Sinha [11] as successive $L^{2}$ projections, laying the foundation for geometric corrections making use of higher derivatives of the square root embedding and curvature information “beyond a certain order" (cf. Theorem 5).
  • These basic results, and those of Bhattacharyya/Pillai–Sinha-type, are further strengthened by a careful analysis of the form of the jets $\eta_{k}$ expressed in terms of the raw score and its derivatives $\partial_{\theta}^{k}\log f(\cdot;\theta)$ via the Faà di Bruno formula and the exponential Bell polynomials. Specifically, we make use of a curvature-aware term, akin to that in the CRB case in Theorem 2, along with orthogonal components arising from Faà di Bruno expansions of the embedding; this result can be found in Proposition 8.
  • The proposed framework is exemplified by applying it to explicit models, demonstrating strict (and fully analytical) improvement over both classical CRBs and algebraic higher-order bounds.
To sum up, this work unifies algebraic and geometric approaches to variance lower bounds and represents, to the best of our knowledge, the first use of extrinsic curvature to sharpen the CRB. It opens new avenues in the intersection of information geometry, Hilbert space theory, and statistical estimation.
2 Preliminaries
In this section, we fix notation and the geometric framework, including square-root embedding, connections, and second fundamental form, used in the sequel. The section is concluded with a brief sketch of our plan to use these ideas to improve upon the CRB and some higher-order versions of the kind worked out in [8, 11].
2.1 Notation, square-root embedding, and assumptions
Let $\mathcal{P}=\{P_{\theta}:\theta\in\Theta\subset\mathbb{R}\}$ be a smooth parametric family of probability measures on a measurable space such that each $P_{\theta}$ has a strictly positive density $f(\cdot;\theta)$ with respect to a fixed base measure $\mu$.
We work with a statistical manifold embedded in the fixed Hilbert space $L^{2}(\mu)$ via the square root map. Consider the square root embedding
$s:\Theta\to L^{2}(\mu),\qquad s(\theta):=s_{\theta}=\sqrt{f(\cdot;\theta)}.$
Denote the $L^{2}(\mu)$ Hilbert space inner product, norm, and orthogonal projection onto a subspace $V$ by
$\langle u,v\rangle:=\int u(x)v(x)\,d\mu(x),\qquad\|u\|^{2}=\langle u,u\rangle,\text{ and }\mathrm{Proj}_{V}(\cdot).$
We have $s_{\theta}\in L^{2}(\mu)$ and $\|s_{\theta}\|^{2}=1$ for all $\theta$. Expectation under $P_{\theta}$ is denoted $\mathbb{E}_{\theta}[\cdot]$. For any measurable $h$, we have $\mathbb{E}_{\theta}[h(X)]=\int h(x)f(x;\theta)\,d\mu(x)=\langle h,s_{\theta}^{2}\rangle$.
Assumption 1.
Assume
$\eta_{k}(\theta):=\partial_{\theta}^{k}s_{\theta},\qquad k\geq 1,$
exist and are in $L^{2}(\mu)$. Further, these are assumed to be linearly independent. Differentiation under the integral sign is permitted; with $T(\cdot):\mathbb{E}_{\theta}[T]<\infty$,
$\frac{\partial}{\partial\theta}\int T(x)f(x;\theta)d\mu=\int T(x)\frac{\partial f(x;\theta)}{\partial\theta}d\mu,$
whenever the right hand side is finite.
Define the (raw) score and higher derivatives of the log-density:
$Y_{k}(x;\theta):=\partial_{\theta}^{k}\log f(x;\theta),\qquad k\geq 1,$
and the Fisher information $\mathcal{I}(\theta):=\mathbb{E}_{\theta}[Y_{1}^{2}]$.
We have, in particular,
$\eta_{1}(\theta)=\partial_{\theta}s_{\theta}=\tfrac{1}{2}Y_{1}\,s_{\theta},$
and
$\langle\eta_{1},\eta_{1}\rangle=\tfrac{1}{4}\mathbb{E}_{\theta}[Y_{1}^{2}]=\tfrac{1}{4}\mathcal{I}(\theta).$
The Fisher–Rao metric on the 1-dimensional parameter manifold is $g(\partial_{\theta},\partial_{\theta})=\mathcal{I}(\theta)$, with $\partial_{\theta}$ denoting the coordinate vector field. Equip $s(\Theta)$ with the induced metric from $L^{2}(\mu)$. The pullback metric from the square-root embedding then equals $\tfrac{1}{4}\mathcal{I}(\theta)$.
Fix $m\geq 1$. Use Assumption 1 to define the finite-dimensional jet subspace
$\mathcal{T}_{m}(\theta):=\operatorname{span}\{\eta_{1}(\theta),\eta_{2}(\theta),\dots,\eta_{m}(\theta)\}\subset L^{2}(\mu),$
and the Gram matrix
$G(\theta)=(G_{ij}(\theta))_{1\leq i,j\leq m},\qquad G_{ij}(\theta):=\langle\eta_{i},\eta_{j}\rangle.$ (1)
By definition and assumed linear independence, the matrix $G(\theta)$ is symmetric positive definite.
2.2 Geometric preliminaries
We now proceed to considering some fundamental geometric and miscellaneous preliminaries (the latter in the next subsection) that serve to make our presentation self-contained.
Ambient (flat) derivative and induced connection
As alluded to, we view $L^{2}(\mu)$ as the ambient Hilbert space with the standard flat metric connection $\nabla$; the ambient directional derivative in the coordinate direction $\partial_{\theta}$ (that is, the pushforward of the coordinate field) acts in the natural way:
$\nabla_{\partial_{\theta}}\eta_{j}=\partial_{\theta}\eta_{j}=\eta_{j+1},\qquad j\geq 1,$
We trivially have that $\nabla$ satisfies metric-compatibility and torsion-freeness:
$X\langle U,V\rangle=\langle\nabla_{X}U,V\rangle+\langle U,\nabla_{X}V\rangle,\qquad\nabla_{X}Y-\nabla_{Y}X=[X,Y],$
for arbitrary smooth ambient vector fields $U,V,X,Y$.
Consider the induced connection obtained by orthogonally projecting the ambient derivative onto $\mathcal{T}_{m}$:
$\nabla^{\mathrm{ind}}_{\partial_{\theta}}\eta_{m}:=\operatorname{Proj}_{\mathcal{T}_{m}}(\eta_{m+1}).$ (2)
Write the induced connection coefficients in the basis $\{\eta_{k}\}_{k=1}^{m}$:
$\nabla^{\mathrm{ind}}_{\partial_{\theta}}\eta_{m}=\sum_{k=1}^{m}\Gamma^{\mathrm{proj}}_{k}\,\eta_{k},$
where we have suppressed the dependence of the row vector $\Gamma^{\mathrm{proj}}$, with components $\Gamma^{\mathrm{proj}}_{k}$, on $m$ (and ignored the risk of confusion!). With a similar notational ambiguity, define the row vector $v(\theta)$ by
$v_{j}(\theta):=\langle\eta_{m+1},\eta_{j}\rangle,\qquad 1\leq j\leq m.$ (3)
Solving the normal equations for the orthogonal projection yields the compact formula
$\Gamma^{\mathrm{proj}}(\theta)=v(\theta)\,G(\theta)^{-1}.$ (4)
In particular, for $m=1$ we have the identity
$\Gamma^{\mathrm{proj}}=\frac{\langle\eta_{2},\eta_{1}\rangle}{\langle\eta_{1},\eta_{1}\rangle}.$ (5)
Again with $m=1$, it is standard to show that the induced connection $\nabla^{\mathrm{ind}}$ on the tangent space $Ts(\Theta)$ obtained by orthogonally projecting the ambient derivative
[…]
where $\widetilde{Y}$ is any ambient extension of the vector field $Y$, is well-defined; that is, it is independent of extension. It is also standard that it is torsion-free and metric-compatible with respect to the pull back metric, and hence coincides with the Levi–Civita connection of the immersed statistical manifold $s(\Theta)$. We refer interested readers to standard references on Riemannian geometry such as [22]. The motivation behind (2) for $m>1$ should also now be clear.
Remark 1.
Note that we work with the fixed Hilbert space $L^{2}(\mu)$ because it provides a fixed inner product and a metric-compatible Levi–Civita connection. The ambient connection in $L^{2}(P_{\theta})$ is not compatible with the metric there. Of course, one could enforce metric-compatibility by hand (by “redefining the Christoffel symbols”), but then we are no longer working with the Levi–Civita connection of the ambient Hilbert space – the geometry is compromised – and we would be simply importing back the $L^{2}(\mu)$ structure in disguise.
Second fundamental form (vectors) and normal bundle
Define the “second fundamental vector" by taking the normal component of the ambient derivative:
$\mathrm{\Pi}_{m}\equiv\mathrm{II}(\eta_{1},\eta_{m}):=\eta_{m+1}-\sum_{k=1}^{m}\Gamma^{\mathrm{proj}}_{k}\,\eta_{k}.$ (6)
For $m=1$, this is just the standard second fundamental form. By construction, $\mathrm{II}(\eta_{1},\eta_{m})\in\mathcal{T}_{m}(\theta)^{\perp}$, i.e.
$\langle\mathrm{II}(\eta_{1},\eta_{m}),\eta_{i}\rangle=0\qquad\text{for }1\leq i\leq m.$
Define the curvature (normal) span
$\mathcal{N}_{m}(\theta):=\operatorname{span}\{\mathrm{II}_{m}\}\subset\mathcal{T}_{m}(\theta)^{\perp}.$ (7)
2.3 Faà di Bruno and exponential Bell polynomials
The combinatorics that repeatedly appear in our derivations, when expressing the jets $\eta_{j}$ in terms of the raw scores $Y_{j}$, are precisely encoded by the Faà di Bruno formula and the exponential Bell polynomials. Below we state the identities in the form used in the paper and give a few expressions that result for relevant quantities such as $v,G$, which in turn determine the induced connection via $\Gamma^{\text{proj}}$, in terms of (mixed) moments of the $Y_{j}$.
Recall $s_{\theta}(x)=\sqrt{f(x;\theta)}$. For brevity, we suppress the dependence on $x$ and $\theta$ when no ambiguity arises. Set
$u(\theta):=\tfrac{1}{2}\log f(\,\cdot\,;\theta),\qquad\text{so that}\qquad s_{\theta}=\exp\big(u(\theta)\big).$
Write $u^{(r)}:=\partial_{\theta}^{r}u(\theta)$. Recall the (first-order) score $Y_{1}=\partial_{\theta}\log f$ so that
$u^{(1)}=\tfrac{1}{2}Y_{1},\quad u^{(2)}=\tfrac{1}{2}Y_{2},\ \dots,\ u^{(k)}=\tfrac{1}{2}Y_{k}.$
Exponential (complete) Bell polynomials.
The complete exponential Bell polynomial $B_{k}(x_{1},\dots,x_{k})$ is defined for $k\geq 1$ by
$B_{k}(x_{1},\dots,x_{k})\;=\;\sum_{m=1}^{k}B_{k,m}(x_{1},\dots,x_{k-m+1}),$
where $B_{k,m}$ are the partial Bell polynomials. The explicit combinatorial formula for the partial Bell polynomials is
$B_{k,m}(x_{1},\dots,x_{k-m+1})\;=\;\sum_{\begin{subarray}{c}(j_{1},\dots,j_{k-m+1})\\[2.0pt] j_{i}\geq 0\end{subarray}}\frac{k!}{j_{1}!j_{2}!\cdots j_{k-m+1}!}\;\prod_{i=1}^{\,k-m+1}\left(\frac{x_{i}}{i!}\right)^{j_{i}},$ (8)
the sum taken over all nonnegative integer solutions $(j_{1},\dots,j_{k-m+1})$ of the two constraints
$\sum_{i=1}^{k-m+1}j_{i}\;=\;m,\qquad\sum_{i=1}^{k-m+1}i\,j_{i}\;=\;k.$
Intuitively, each term of $B_{k,m}$ corresponds to a partition of the set $\{1,\dots,k\}$ into $m$ blocks, where $j_{i}$ counts the number of blocks of size $i$.
Faà di Bruno (exponential form).
A convenient form of Faà di Bruno for the $k$-th derivative of an exponential composition $s(\theta)=\exp(u(\theta))$ is
$\frac{d^{k}}{d\theta^{k}}\,e^{u(\theta)}\;=\;e^{u(\theta)}\,B_{k}\!\big(u^{(1)}(\theta),u^{(2)}(\theta),\dots,u^{(k)}(\theta)\big).$ (9)
Equivalently, with the partial Bell polynomials,
$\frac{d^{k}}{d\theta^{k}}\,e^{u(\theta)}\;=\;e^{u(\theta)}\sum_{m=1}^{k}B_{k,m}\!\big(u^{(1)},u^{(2)},\dots,u^{(k-m+1)}\big).$
This identity is the standard Faà di Bruno formula specialized to the case when the outer function in the composition is the exponential function.
Application to the square-root embedding.
Apply (9) with $u(\theta)=\tfrac{1}{2}\log f(\cdot;\theta)$. Since $s_{\theta}=\exp(u(\theta))$, we obtain for every integer $k\geq 1$:
$\eta_{k}(\cdot\,;\theta)\;=\;\partial_{\theta}^{k}s_{\theta}\;=\;s_{\theta}\;B_{k}\!\big(u^{(1)},u^{(2)},\dots,u^{(k)}\big)\;=\;s_{\theta}\;B_{k}\!\Big(\tfrac{1}{2}Y_{1},\tfrac{1}{2}Y_{2},\dots,\tfrac{1}{2}Y_{k}\Big).$ (10)
This is the compact identity of great relevance in the paper. The above expresses each $k$-th jet $\eta_{k}$ as the product of the base function $s_{\theta}$ and a finite linear combination of monomials of the form $Y_{i_{1}}Y_{i_{2}}\cdots Y_{i_{k}}$, with coefficients determined by the Bell polynomials.
Computation of the Gram matrix $G$ and vector $v$.
Using (10) the Gram matrix entries are
$G_{ij}=\langle\eta_{i},\eta_{j}\rangle=\int s_{\theta}^{2}\;B_{i}\!\Bigl(\tfrac{1}{2}Y_{1},\dots,\tfrac{1}{2}Y_{i}\Bigr)\;B_{j}\!\Bigl(\tfrac{1}{2}Y_{1},\dots,\tfrac{1}{2}Y_{j}\Bigr)\,d\mu,$
which is equivalently an expectation under $P_{\theta}$:
$\quad G_{ij}\;=\;\mathbb{E}_{\theta}\!\Big[B_{i}\bigl(\tfrac{1}{2}Y_{1},\dots,\tfrac{1}{2}Y_{i}\bigr)\;B_{j}\bigl(\tfrac{1}{2}Y_{1},\dots,\tfrac{1}{2}Y_{j}\bigr)\Big]\quad.$
Similarly, the vector $v$ defined in (3) has entries
$v_{j}=\langle\eta_{m+1},\eta_{j}\rangle=\mathbb{E}_{\theta}\!\Big[B_{m+1}\bigl(\tfrac{1}{2}Y_{1},\dots,\tfrac{1}{2}Y_{m+1}\bigr)\;B_{j}\bigl(\tfrac{1}{2}Y_{1},\dots,\tfrac{1}{2}Y_{j}\bigr)\Big].$
Both $G$ and $v$ are therefore computable from the joint moments of the vector $(Y_{1},\dots,Y_{m+1})$.
Lower-order expansions.
For reader convenience, we record the first few cases obtained from (10) by using the definitions $B_{1}(x_{1})=x_{1}$, $B_{2}(x_{1},x_{2})=x_{1}^{2}+x_{2}$, $B_{3}(x_{1},x_{2},x_{3})=x_{1}^{3}+3x_{1}x_{2}+x_{3}$:
$\eta_{1}$ $=s_{\theta}\;\tfrac{1}{2}Y_{1},$ (11)
$\eta_{2}$ $=s_{\theta}\;\Big(\tfrac{1}{4}Y_{1}^{2}+\tfrac{1}{2}Y_{2}\Big),$ (12)
$\eta_{3}$ $=s_{\theta}\;\Big(\tfrac{1}{8}Y_{1}^{3}+\tfrac{3}{4}Y_{1}Y_{2}+\tfrac{1}{2}Y_{3}\Big).$ (13)
With this, we may compute the entries of $G$ and $v$, and the resulting $\mathrm{II}_{m}$.
Order 1: In the scalar $m=1$ case, from (11),
$G_{11}=\langle\eta_{1},\eta_{1}\rangle=\tfrac{1}{4}\mathbb{E}_{\theta}[Y_{1}^{2}].$
$v_{1}=\langle\eta_{2},\eta_{1}\rangle=\Big\langle\bigl(\tfrac{1}{2}Y_{2}+\tfrac{1}{4}Y_{1}^{2}\bigr)s_{\theta},\tfrac{1}{2}Y_{1}s_{\theta}\Big\rangle=\tfrac{1}{4}\mathbb{E}_{\theta}[Y_{2}Y_{1}]+\tfrac{1}{8}\mathbb{E}_{\theta}[Y_{1}^{3}].$
Further,
$\Gamma^{\mathrm{proj}}_{1}=\frac{v_{1}}{G_{11}}.$
Therefore the second fundamental form equals
$\mathrm{II}_{1}=\eta_{2}-\Gamma^{\mathrm{proj}}_{1}\,\eta_{1}=s_{\theta}\Big(\tfrac{1}{2}Y_{2}+\tfrac{1}{4}Y_{1}^{2}-\frac{v_{1}}{G_{11}}\cdot\tfrac{1}{2}Y_{1}\Big).$ (14)
Order 2: We now rely also on (12) and (13). The relevant Gram entries are
$G_{12}$ $=\langle\eta_{1},\eta_{2}\rangle=\tfrac{1}{4}\mathbb{E}_{\theta}[Y_{1}Y_{2}]+\tfrac{1}{8}\mathbb{E}_{\theta}[Y_{1}^{3}],$
$G_{22}$ $=\langle\eta_{2},\eta_{2}\rangle=\mathbb{E}_{\theta}\Big[\bigl(\tfrac{1}{2}Y_{2}+\tfrac{1}{4}Y_{1}^{2}\bigr)^{2}\Big]=\tfrac{1}{4}\mathbb{E}_{\theta}[Y_{2}^{2}]+\tfrac{1}{4}\mathbb{E}_{\theta}[Y_{2}Y_{1}^{2}]+\tfrac{1}{16}\mathbb{E}_{\theta}[Y_{1}^{4}].$
Similarly,
$v_{1}=\langle\eta_{3},\eta_{1}\rangle=\tfrac{1}{4}\mathbb{E}_{\theta}[Y_{3}Y_{1}]+\tfrac{3}{8}\mathbb{E}_{\theta}[Y_{1}^{2}Y_{2}]+\tfrac{1}{16}\mathbb{E}_{\theta}[Y_{1}^{4}],$
and
$v_{2}=\langle\eta_{3},\eta_{2}\rangle=\mathbb{E}_{\theta}\Big[\bigl(\tfrac{1}{2}Y_{3}+\tfrac{3}{4}Y_{1}Y_{2}+\tfrac{1}{8}Y_{1}^{3}\bigr)\bigl(\tfrac{1}{2}Y_{2}+\tfrac{1}{4}Y_{1}^{2}\bigr)\Big],$
which also expands into a linear combination of mixed moments.
With the $2\times 2$ matrix $G=[G_{ij}]$ and $1\times 2$ vector $v=[v_{j}]$, we may then compute the row vector $\Gamma^{\mathrm{proj}}=vG^{-1}$. Subsequently, using
$\mathrm{II}_{2}=\eta_{3}-\Gamma^{\mathrm{proj}}_{1}\,\eta_{1}-\Gamma^{\mathrm{proj}}_{2}\,\eta_{2},$
and substituting the exact Bell polynomial-based identities for $\eta_{1},\eta_{2},\eta_{3}$ gives the fully explicit expansion of $\mathrm{II}_{2}$.
2.4 Outlook and road ahead
Consider any unbiased estimator $T(X)$ ($X\sim P_{\theta}$) of the scalar parameter $\theta$ and define the centered estimation error $Z_{0}:=T(X)-\theta$ viewed as an element of $L^{2}(P_{\theta})$ or, equivalently, focus on the vector $\widetilde{Z}_{0}:=Z_{0}s_{\theta}\in L^{2}(\mu)$. Similarly, also define $\widetilde{Y}_{1}:=Y_{1}s_{\theta},\widetilde{Y}_{2}:=Y_{2}s_{\theta},\dots,\widetilde{Y}_{m}:=Y_{m}s_{\theta}$ and $\widetilde{\mathcal{T}}_{m}:=\mathrm{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{m}\}.$
Now, it is easy enough to show (we do it below) that $\|\mathrm{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}\|^{2}=\|\mathrm{Proj}_{\widetilde{\mathcal{T}}_{1}}\widetilde{Z}_{0}\|^{2}=\tfrac{1}{\mathcal{I}(\theta)}$, which lower bounds the variance $\mathrm{Var}_{\theta}[T]$. Our insight is so simple, hence elegant: Orthogonally decompose
$\widetilde{Z}_{0}=\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}+R_{m},\qquad R_{m}\in\mathcal{T}_{m}^{\perp}.$
Then the squared norm $\|\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\|^{2}$ is the projection-based portion of the variance (for $m=1$ it reproduces the CRB, as just mentioned); the residual $\|R_{m}\|^{2}$ is the unexplained variance by the $m$-term projection. Projecting $R_{m}$ further onto the curvature span $\mathcal{N}_{m}$ (see (7)) yields an explicit curvature correction. In summary, we may better the CRB by considering the curvature-induced residual for $m=1$. We emphasize again that this idea is distinct from the method of Bhattacharyya [8] which uses higher-order derivatives of the score and does not exploit the extrinsic geometry, as we do, via the square root embedding.
With $m>1$, Bhattacharyya’s [8, 9, 10]/Pillai and Sinha’s result [11] can be considered as improving the CRB using the subspace $\widetilde{\mathcal{T}}_{m}$, and not $\mathcal{T}_{m}=\mathrm{span}\{\eta_{1},\dots,\eta_{m}\}$, in the projection above (strictly speaking, Bhattacharyya/Pillai–Sinha consider the ratios of the derivatives of the density $f(x;\theta)$ to the density function, and not directly the derivatives of the log-likelihood as already mentioned). In any case, we prefer to work with $\mathcal{T}_{m}$ in developing our preliminary curvature-correction result (cf. Theorem 5 below) for reasons related to those quoted in Remark 1; in short, this approach offers an intuitive and correct geometric picture. One may ask: How does the subspace $\mathcal{T}_{m}$ compare against $\widetilde{\mathcal{T}}_{m}$ or, equivalently, how does our theorem relate to those of the nature of Bhattacharyya/Pillai–Sinha? While the spaces $\widetilde{\mathcal{T}}_{1}=\mathcal{T}_{1}$, the answer is a bit more subtle in the $m>1$ case, and we refer interested readers to Section 4.3 for some stronger (in terms of improvement in variance lower bounds) results that shed light on this scenario and Section 4.4 for further discussion. Our strongest result for $m>1$, which betters bounds of the kind obtained by Bhattacharyya/Pillai and Sinha, appears in Proposition 8.
3 Curvature-corrected variance bounds – CRB case
We now state and prove the main bounds where, unless mentioned otherwise and as has been the case throughout, all inner products are those of the fixed Hilbert space $L^{2}(\mu)$, and all objects depend on the true parameter value $\theta$ (but we suppress the explicit $\theta$ argument, as has been frequent practice, for notational brevity). Recall that we have established (cf. (5) or (14)) that the second fundamental form is the normal component of $\eta_{2}$:
$\mathrm{II}(\eta_{1},\eta_{1})\;=\;\eta_{2}-\frac{\langle\eta_{2},\eta_{1}\rangle}{\langle\eta_{1},\eta_{1}\rangle}\,\eta_{1}\;\in\;\mathcal{T}_{1}^{\perp}.$
The proof of the $m>1$ case in Theorem 5 essentially covers that of $m=1$ as well, but we state the $m=1$ case separately since it corresponds to improving the classical CRB.
Theorem 2 (Residual variance and CRB improvement, $m=1$).
Assume the model satisfies Assumption 1. Fix the true parameter value $\theta$ and let $T(X)$ be an estimator of $\theta$ which is unbiased (in a neighbourhood of $\theta$). Define the centered error
$Z_{0}(X):=T(X)-\theta.$
With $s_{\theta}=\sqrt{f(\cdot;\theta)}\in L^{2}(\mu)$, consider the ambient error vector, the one-dimensional tangent space, and residual, respectively:
$\widetilde{Z}_{0}:=Z_{0}\,s_{\theta},\quad\mathcal{T}_{1}=\operatorname{span}\{\eta_{1}\}\subset L^{2}(\mu),\quad R_{1}:=\widetilde{Z}_{0}-\operatorname{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}\in\mathcal{T}_{1}^{\perp}.$
Finally denote the Fisher information by $\mathcal{I}(\theta)=\mathbb{E}_{\theta}[Y_{1}^{2}]$, where the raw score $Y_{1}=\partial_{\theta}\log f(X;\theta)$.
Then the following statements hold.
  1. 1.
    (Variance decomposition/CRB recovery)
    $\operatorname{Var}_{\theta}[T]\;=\;\big\|\operatorname{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}\big\|^{2}\;+\;\|R_{1}\|^{2}\;\geq\;\frac{1}{\mathcal{I}(\theta)}.$
    Specifically, under the stated unbiasedness and regularity assumptions
    $\big\|\operatorname{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}\big\|^{2}\;=\;\frac{1}{\mathcal{I}(\theta)},$
    so the classical Cramér–Rao lower bound is obtained from the $m=1$ projection of the centered error onto the tangent space $\mathcal{T}_{1}$.
  2. 2.
    (Curvature lower bound)
    $\|R_{1}\|^{2}\;\geq\;\frac{\big\langle\widetilde{Z}_{0},\;\mathrm{II}(\eta_{1},\eta_{1})\big\rangle^{2}}{\big\|\mathrm{II}(\eta_{1},\eta_{1})\big\|^{2}},$ (15)
    with equality if and only if $R_{1}$ is collinear (in $L^{2}(\mu)$) with $\mathrm{II}(\eta_{1},\eta_{1})$.
In particular, combining our conclusions gives the curvature-corrected bound
$\operatorname{Var}_{\theta}[T]\;=\;\|\widetilde{Z}_{0}\|^{2}\;\geq\;\frac{1}{\mathcal{I}(\theta)}\;+\;\frac{\big\langle\widetilde{Z}_{0},\;\mathrm{II}(\eta_{1},\eta_{1})\big\rangle^{2}}{\big\|\mathrm{II}(\eta_{1},\eta_{1})\big\|^{2}}.$
Proof.
1. Orthogonally decompose the ambient vector $\widetilde{Z}_{0}$ into tangential and normal parts relative to $\mathcal{T}_{1}$:
$\widetilde{Z}_{0}=\operatorname{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}+R_{1},\qquad\operatorname{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}\in\mathcal{T}_{1},\ R_{1}\in\mathcal{T}_{1}^{\perp}.$
Noting that $\|\widetilde{Z}_{0}\|^{2}=\mathbb{E}_{\theta}[(T-\theta)^{2}]=\operatorname{Var}_{\theta}[T]$, taking squared norms, and using orthogonality yields the variance decomposition
$\operatorname{Var}_{\theta}[T]=\big\|\operatorname{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}\big\|^{2}+\|R_{1}\|^{2}.$
The orthogonal projection onto the one-dimensional subspace $\mathcal{T}_{1}=\operatorname{span}\{\eta_{1}\}$ has squared norm
$\big\|\operatorname{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}\big\|^{2}=\frac{\langle\widetilde{Z}_{0},\eta_{1}\rangle^{2}}{\langle\eta_{1},\eta_{1}\rangle}.$
Using the identity $\eta_{1}=\tfrac{1}{2}Y_{1}s_{\theta}$ we compute
$\langle\eta_{1},\eta_{1}\rangle=\frac{1}{4}\,\mathbb{E}_{\theta}[Y_{1}^{2}]=\frac{1}{4}\,\mathcal{I}(\theta),\qquad\langle\widetilde{Z}_{0},\eta_{1}\rangle=\big\langle Z_{0}s_{\theta},\tfrac{1}{2}Y_{1}s_{\theta}\big\rangle=\tfrac{1}{2}\,\mathbb{E}_{\theta}[(T-\theta)Y_{1}].$
Under the unbiasedness and regularity hypotheses we may differentiate $\mathbb{E}_{\theta}[T]=\theta$ with respect to $\theta$ and interchange derivative and expectation; this yields $\mathbb{E}_{\theta}[TY_{1}]=1$. Similarly, since $f(\cdot;\theta)$ integrates to $1$, we get $\mathbb{E}_{\theta}[Y_{1}]=0$ and hence $\mathbb{E}_{\theta}[(T-\theta)Y_{1}]=1$. Therefore
$\langle\widetilde{Z}_{0},\eta_{1}\rangle=\tfrac{1}{2},\quad\text{and}\quad\big\|\operatorname{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}\big\|^{2}=\frac{(1/2)^{2}}{\mathcal{I}(\theta)/4}=\frac{1}{\mathcal{I}(\theta)}.$
2. Since $\mathrm{II}(\eta_{1},\eta_{1})\in\mathcal{T}_{1}^{\perp}$ and $R_{1}\in\mathcal{T}_{1}^{\perp}$, both vectors lie in the same subspace $\mathcal{T}_{1}^{\perp}$. In that subspace, the elementary one-dimensional projection inequality gives
$\|R_{1}\|^{2}\;\geq\;\frac{\langle R_{1},\mathrm{II}(\eta_{1},\eta_{1})\rangle^{2}}{\|\mathrm{II}(\eta_{1},\eta_{1})\|^{2}},$
with equality iff $R_{1}$ is collinear with $\mathrm{II}(\eta_{1},\eta_{1})$. Finally, because $\operatorname{Proj}_{\mathcal{T}_{1}}\widetilde{Z}_{0}$ is orthogonal to $\mathrm{II}(\eta_{1},\eta_{1})$, we have
$\langle R_{1},\mathrm{II}(\eta_{1},\eta_{1})\rangle=\langle\widetilde{Z}_{0},\mathrm{II}(\eta_{1},\eta_{1})\rangle,$
which yields (15) and completes the proof. ∎
4 Higher-order CRB and curvature corrections in the square-root embedding
Needless to (re)state, we keep Assumption 1 so that all quantities below are well defined in the fixed Hilbert space $L^{2}(\mu)$. We first state the Hilbert-space formulation of the Bhattacharyya bound in $L^{2}(P_{\theta})$ as conceptualized in [11] with the caveat that we use derivatives of the log likelihood and not the ratios $\frac{\partial_{\theta}^{k}f}{f}$, show its equivalent formulation in the ambient space $L^{2}(\mu)$, and then present a basic curvature-corrected bound formulated relative to the jet subspace $\mathcal{T}_{m}=\operatorname{span}\{\eta_{1},\dots,\eta_{m}\}$; the core idea here is the same as for the $m=1$ case. This bound is finally strengthened in Proposition 8 in Section 4.3, where it is explicitly shown that we improve over the Bhattacharyya-type bounds in the framework of Pillai and Sinha using derivatives of the raw score.
4.1 Bhattacharyya/Pillai–Sinha-type bounds in $L^{2}(P_{\theta})$ and its square-root representation
The following theorem formalizes, strengthens, and rephrases the core insight of Pillai and Sinha [11], who construct a sequence of variance lower bounds using higher-order score functions systematically covering the Bhattacharyya variants [8]. We claim no real novelty in the following result, but include the same anyway for the following reasons: While their work asserts that equality in the $m$-th-order bound results if estimator error lies in the span of score functions up to order $m$, our formulation provides a characterization of equivalence using Hilbert space projections, and sets the stage for further improvements to be derived later via basic Riemannian geometric arguments.
As regards notation, consider $L^{2}(P_{\theta})$ with inner product $\langle f,g\rangle_{\theta}=\mathbb{E}_{\theta}[f(X)g(X)]$ and norm $\|f\|_{\theta}^{2}=\langle f,f\rangle_{\theta}$. We use $\operatorname{Proj}^{\theta}_{V}(\cdot)$ to denote orthogonal projection to a subspace $V$ of $L^{2}(P_{\theta})$.
Theorem 3 (Hilbert space characterization of higher-order CRB [8, 11]).
Let $T(X)$ be an unbiased estimator of a scalar parameter $\theta$, and define the centered error $Z_{0}:=T(X)-\theta\in L^{2}(P_{\theta})$. For each $k\geq 1$, consider the $k$-th score function $Y_{k}=\partial_{\theta}^{k}\log f(X;\theta)$ ($\in L^{2}(P_{\theta})$), and define the finite-dimensional subspace:
$\mathcal{T}^{Y}_{m}:=\operatorname{span}\{Y_{1},\dots,Y_{m}\}\subset L^{2}(P_{\theta}),$
where the $Y_{i}$ are assumed to be linearly independent. Let $G^{Y}\in\mathbb{R}^{m\times m}$ be the Gram matrix with entries $G^{Y}_{ij}:=\langle Y_{i},Y_{j}\rangle_{\theta}$, and define $b\in\mathbb{R}^{m}$ by $b_{i}:=\langle Z_{0},Y_{i}\rangle_{\theta}$. Denote the variance $\sigma^{2}:=\|Z_{0}\|_{\theta}^{2}=\operatorname{Var}_{\theta}[T(X)]$, and define the matrix:
$B^{(m)}:=\mathbb{E}_{\theta}\left[\begin{bmatrix}Z_{0}\\ Y_{1}\\ \vdots\\ Y_{m}\end{bmatrix}\begin{bmatrix}Z_{0}\\ Y_{1}\\ \vdots\\ Y_{m}\end{bmatrix}^{\top}\right]=\begin{bmatrix}\sigma^{2}&b^{\top}\\ b&G^{Y}\end{bmatrix}$.
Then we have $\operatorname{Var}_{\theta}[T(X)]\geq b^{\top}(G^{Y})^{-1}b=\|\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{k}}(Z_{0})\|_{\theta}^{2}$. Further, the following are equivalent:
  1. 1.
    $Z_{0}\in\mathcal{T}^{Y}_{m}$, i.e., the projection error vanishes: $\|Z_{0}-\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{m}}(Z_{0})\|_{\theta}^{2}=0$;
  2. 2.
    $B^{(m)}$ is singular; specifically, $\left(\left[(B^{(m)})^{-1}\right]_{11}\right)^{-1}=0$.
Define $\mathcal{B}^{(k)}:=\|\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{k}}(Z_{0})\|_{\theta}^{2}$ for $k\geq 1$. If either of the above equivalent conditions hold, then $\mathcal{B}^{(1)}\leq\mathcal{B}^{(2)}\leq\dots$, and for all $k\geq m$,
$\mathcal{B}^{(k)}=\operatorname{Var}_{\theta}[T(X)].$
Proof.
Since $T(X)$ is unbiased, $Z_{0}\in L^{2}(P_{\theta})$ satisfies $\mathbb{E}_{\theta}[Z_{0}]=0$ and its variance equals $\|Z_{0}\|_{\theta}^{2}$.
We can write:
$\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{m}}(Z_{0})=\sum_{k=1}^{m}a_{k}Y_{k},\quad\text{where }a=(G^{Y})^{-1}b,$
and the squared projection norm is:
$\|\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{m}}(Z_{0})\|_{\theta}^{2}=b^{\top}(G^{Y})^{-1}b.$
Thus, the residual variance is:
$0\leq\|Z_{0}-\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{m}}(Z_{0})\|_{\theta}^{2}=\|Z_{0}\|_{\theta}^{2}-b^{\top}(G^{Y})^{-1}b=\sigma^{2}-b^{\top}(G^{Y})^{-1}b.$
Now recall that the inverse of the block matrix $B^{(m)}$ has:
$\left[(B^{(m)})^{-1}\right]_{11}=\frac{1}{\sigma^{2}-b^{\top}(G^{Y})^{-1}b}=\frac{1}{\|Z_{0}-\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{m}}(Z_{0})\|_{\theta}^{2}},$
from the Schur complement of $G^{Y}$ in $B^{(m)}$. Therefore, the following are equivalent:
(i) $Z_{0}\in\mathcal{T}^{Y}_{m}$ or $\|Z_{0}-\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{m}}(Z_{0})\|_{\theta}^{2}=0$,
(ii) $B^{(m)}$ is singular and $\left(\left[(B^{(m)})^{-1}\right]_{11}\right)^{-1}=0$.
Since $Z_{0}\in\mathcal{T}^{Y}_{m}$, projections onto a larger space, which are nondecreasing in norm until $k<m$, stabilize at $k=m$; i.e. for $k>m$, since $\mathcal{T}^{Y}_{k}\supseteq\mathcal{T}^{Y}_{m}$, projections give the same result. Thus, $\|\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{k}}(Z_{0})\|_{\theta}^{2}=\|Z_{0}\|_{\theta}^{2}$ for all $k\geq m$. ∎
Remark 2.
It is obvious, again for $m=1$, that the general bound $b^{\top}(G^{Y})^{-1}b$ becomes:
$\mathcal{B}^{(1)}=\frac{1^{2}}{\mathcal{I}(\theta)}=\frac{1}{\mathcal{I}(\theta)}.$
Lemma 4 (Square-root isometry and score-image representation).
Let $s_{\theta}=\sqrt{f(\cdot;\theta)}$ and define the map $\mathcal{J}:L^{2}(P_{\theta})\to L^{2}(\mu)$ by $(\mathcal{J}h)(x):=h(x)\,s_{\theta}(x)$. For $k\geq 1$, define the score-image $\widetilde{Y}_{k}:=\mathcal{J}(Y_{k})=Y_{k}s_{\theta}$ and $\widetilde{Z}_{0}:=\mathcal{J}(Z_{0})=Z_{0}s_{\theta}$. Then
$\langle Y_{i},Y_{j}\rangle_{\theta}=\langle\widetilde{Y}_{i},\widetilde{Y}_{j}\rangle,\qquad\langle Z_{0},Y_{i}\rangle_{\theta}=\langle\widetilde{Z}_{0},\widetilde{Y}_{i}\rangle.$
Consequently Theorem 3 is equivalent to the identical projection statement in the fixed Hilbert space $L^{2}(\mu)$ with subspace $\widetilde{\mathcal{T}}_{m}:=\operatorname{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{m}\}$.
Specifically, $\mathcal{B}^{(k)}=\|\operatorname{Proj}^{\theta}_{\mathcal{T}^{Y}_{k}}(Z_{0})\|_{\theta}^{2}=\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{k}}(\widetilde{Z}_{0})\|^{2}$, for $k\geq 1$.
Proof.
The verification is immediate:
$\langle\mathcal{J}h_{1},\mathcal{J}h_{2}\rangle=\int h_{1}(x)h_{2}(x)s_{\theta}(x)^{2}\,d\mu(x)=\int h_{1}(x)h_{2}(x)f(x;\theta)\,d\mu(x)=\langle h_{1},h_{2}\rangle_{\theta}.$
Thus inner products and orthogonal projections are preserved under $\mathcal{J}$, establishing the equivalence. ∎
Example 1 (Saturation at $m=3$ in a symmetric quartic location model).
Consider the location model with density:
$f(x;\theta)=\frac{1}{Z}\exp\left(-(x-\theta)^{4}\right),\quad x\in\mathbb{R}$
where the normalization constant is $Z=\int_{-\infty}^{\infty}\exp(-u^{4})\,du$ via the change of variable $u=x-\theta$. This model is symmetric, smooth, and lies outside the exponential family. We work in the Hilbert space $L^{2}(P_{\theta})$.
Let $X\sim f(x;\theta)$ and define the centered estimator:
$Z_{0}=T(X)-\theta=X-\theta$
The first three (raw) score functions are:
$Y_{1}(x)=\frac{\partial}{\partial\theta}\log f(x;\theta)=4(x-\theta)^{3},\quad Y_{2}(x)=\frac{\partial^{2}}{\partial\theta^{2}}\log f(x;\theta)=-12(x-\theta)^{2}$
$Y_{3}(x)=\frac{\partial^{3}}{\partial\theta^{3}}\log f(x;\theta)=24(x-\theta)$
We compute the bounds:
$\mathcal{B}^{(1)}=\frac{\langle Z_{0},Y_{1}\rangle_{\theta}^{2}}{\langle Y_{1},Y_{1}\rangle_{\theta}},\quad\mathcal{B}^{(2)}=\begin{bmatrix}\langle Z_{0},Y_{1}\rangle_{\theta}&\langle Z_{0},Y_{2}\rangle_{\theta}\end{bmatrix}\cdot\begin{bmatrix}\langle Y_{1},Y_{1}\rangle_{\theta}&\langle Y_{1},Y_{2}\rangle_{\theta}\\ \langle Y_{1},Y_{2}\rangle_{\theta}&\langle Y_{2},Y_{2}\rangle_{\theta}\end{bmatrix}^{-1}\cdot\begin{bmatrix}\langle Z_{0},Y_{1}\rangle_{\theta}\\ \langle Z_{0},Y_{2}\rangle_{\theta}\end{bmatrix}$
$\mathcal{B}^{(3)}=b^{\top}(G^{Y})^{-1}b,\quad\text{with }b=\begin{bmatrix}\langle Z_{0},Y_{1}\rangle_{\theta}\\ \langle Z_{0},Y_{2}\rangle_{\theta}\\ \langle Z_{0},Y_{3}\rangle_{\theta}\end{bmatrix},\quad G^{Y}=\begin{bmatrix}\langle Y_{1},Y_{1}\rangle_{\theta}&\langle Y_{1},Y_{2}\rangle_{\theta}&\langle Y_{1},Y_{3}\rangle_{\theta}\\ \langle Y_{1},Y_{2}\rangle_{\theta}&\langle Y_{2},Y_{2}\rangle_{\theta}&\langle Y_{2},Y_{3}\rangle_{\theta}\\ \langle Y_{1},Y_{3}\rangle_{\theta}&\langle Y_{2},Y_{3}\rangle_{\theta}&\langle Y_{3},Y_{3}\rangle_{\theta}\end{bmatrix}$.
Symmetry-induced orthogonality. Since the model is symmetric around $\theta$, and $Z_{0},Y_{1}$ are odd functions while $Y_{2}$ is even, it follows that:
$\langle Z_{0},Y_{2}\rangle_{\theta}=0,\quad\langle Y_{1},Y_{2}\rangle_{\theta}=0$
Therefore, we get:
$\mathcal{B}^{(2)}=\mathcal{B}^{(1)}<\operatorname{Var}_{\theta}[T]$
Third-order saturation. Since $Y_{3}(x)=24(x-\theta)=24Z_{0}(x)$, we have:
$\langle Z_{0},Y_{3}\rangle_{\theta}=24\cdot\operatorname{Var}_{\theta}[T]$
and $Z_{0}\in\operatorname{span}\{Y_{1},Y_{3}\}$, implying:
$\mathcal{B}^{(3)}=\operatorname{Var}_{\theta}[T]$
Numerical values. Using numerical integration under the base measure $p(u)=\frac{1}{Z}e^{-u^{4}}$, we obtain:
$\mathcal{B}^{(1)}=\mathcal{B}^{(2)}=0.24656,\qquad\mathcal{B}^{(3)}=\operatorname{Var}_{\theta}[T]=0.33799$
This example demonstrates strict improvement of the lower bound at $m=3$ and confirms saturation of the bound sequence at that order.
4.2 Curvature-corrected variance bound, $m>1$
We now state and prove the curvature-corrected bound formulated relative to the jet subspace $\mathcal{T}_{m}$, and not $\widetilde{\mathcal{T}}_{m}$, since this really is the natural generalization of the $m=1$ case (see also Remark 1). Thus we work in the fixed Hilbert space $L^{2}(\mu)$ and the jet subspace $\mathcal{T}_{m}$. The geometric refinement we use is again similar in spirit to the $m=1$ scenario: project the estimator residual onto the jet space, then project the residual onto the curvature span inside the normal bundle.
Theorem 5 (Curvature-corrected variance bound, $m\geq 1$).
Let $T(X)$ be an unbiased estimator of $\theta$ and $Z_{0}=T(X)-\theta$. Define the centered error $\widetilde{Z}_{0}:=Z_{0}s_{\theta}\in L^{2}(\mu)$, where $s_{\theta}=\sqrt{f(\cdot,\theta)}$. For $k\geq 1$, consider $\eta_{k}=\partial_{\theta}^{k}s_{\theta}$ and the jet subspace
$\mathcal{T}_{m}=\operatorname{span}\{\eta_{1},\dots,\eta_{m}\}\subset L^{2}(\mu).$
Consider the second fundamental vector
$\mathrm{II}_{m}:=\mathrm{II}(\eta_{1},\eta_{m})=\eta_{m+1}-\sum_{k=1}^{m}\Gamma^{\mathrm{proj}}_{k}\,\eta_{k},$
where the row vector $\Gamma^{\mathrm{proj}}=vG^{-1}$ with the matrix $G_{ij}=\langle\eta_{i},\eta_{j}\rangle$ and the row vector $v_{j}=\langle\eta_{m+1},\eta_{j}\rangle$. Define the curvature (normal) span
$\mathcal{N}_{m}:=\operatorname{span}\{\mathrm{II}_{m}\}\subset\mathcal{T}_{m}^{\perp}.$
Then the following curvature-corrected variance bound holds:
$\operatorname{Var}_{\theta}[T]\geq\mathcal{B}^{(m)}_{\mathrm{jet}}+\|\operatorname{Proj}_{\mathcal{N}_{m}}(\widetilde{Z}_{0})\|^{2}=:\mathcal{C}^{(m)},$ (16)
where $\mathcal{B}^{(m)}_{\mathrm{jet}}:=\|\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\|^{2}$ is the projection bound relative to the jet subspace. Equality in (16) holds iff $R_{m}:=\widetilde{Z}_{0}-\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\in\mathcal{N}_{m}$.
Proof.
Start with the orthogonal decomposition $\widetilde{Z}_{0}=\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}+R_{m}$, with $R_{m}\in\mathcal{T}_{m}^{\perp}$. The normal space $\mathcal{T}_{m}^{\perp}$ contains the curvature span $\mathcal{N}_{m}$ by construction, so we may orthogonally decompose the residual $R_{m}$ with respect to $\mathcal{N}_{m}$:
$R_{m}=\operatorname{Proj}_{\mathcal{N}_{m}}R_{m}+\bigl(R_{m}-\operatorname{Proj}_{\mathcal{N}_{m}}R_{m}\bigr),$
with the two terms orthogonal. Taking squared norms and summing orthogonal contributions yields
$\|\widetilde{Z}_{0}\|^{2}$
$\|\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\|^{2}+\|\operatorname{Proj}_{\mathcal{N}_{m}}R_{m}\|^{2}+\|R_{m}-\operatorname{Proj}_{\mathcal{N}_{m}}R_{m}\|^{2}\geq\|\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\|^{2}+\|\operatorname{Proj}_{\mathcal{N}_{m}}R_{m}\|^{2},$
which is (16). Note that $\operatorname{Proj}_{\mathcal{N}_{m}}R_{m}=\operatorname{Proj}_{\mathcal{N}_{m}}\widetilde{Z}_{0}$, since $\operatorname{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\perp\mathcal{N}_{m}$. Equality occurs iff the orthogonal remainder $R_{m}-\operatorname{Proj}_{\mathcal{N}_{m}}R_{m}$ vanishes, i.e. $R_{m}\in\mathcal{N}_{m}$. ∎
Remark 3.
There are two finite-dimensional subspaces inside $L^{2}(\mu)$:
$\widetilde{\mathcal{T}}_{k}=\operatorname{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{k}\},\qquad\mathcal{T}_{k}=\operatorname{span}\{\eta_{1},\dots,\eta_{k}\}.$
The Bhattacharyya/Pillai–Sinha-type projection bounds correspond to projecting the estimator residual onto $\widetilde{\mathcal{T}}_{k}$ (equivalently, projecting $Z_{0}$ onto $\mathcal{T}_{k}^{Y}$ in $L^{2}(P_{\theta})$). That projection yields $\mathcal{B}^{(k)}_{\mathrm{score}}=\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{k}}\widetilde{Z}_{0}\|^{2}$.
The curvature correction requires projecting onto the jet subspace $\mathcal{T}_{k}=\operatorname{span}\{\eta_{1},\dots,\eta_{k}\}$, which is the natural subspace to be considered with the induced connection and second fundamental vectors. Even though $\mathcal{T}_{k},\,\widetilde{\mathcal{T}}_{k}$ are both $k$-dimensional subspaces under assumptions of linear independence, in general $\mathcal{T}_{k}\neq\widetilde{\mathcal{T}}_{k}$ for $k\geq 2$, so $\mathcal{B}^{(k)}_{\mathrm{jet}}$ and $\mathcal{B}^{(k)}_{\mathrm{score}}$ are different quantities. Only when the linear spans exactly coincide do the two formulations agree. The one-dimensional spaces $\operatorname{span}\{\widetilde{Y}_{1}\}$ and $\operatorname{span}\{\eta_{1}\}$ coincide because $\eta_{1}=\tfrac{1}{2}\widetilde{Y}_{1}$, which is why the classical CRB is unaffected by the choice. Nevertheless, it is obvious that our result in Theorem 5 offers curvature-aware improvement over $\mathcal{B}^{(m)}_{\mathrm{score}}$ when $\mathcal{B}^{(m)}_{\mathrm{jet}}=\mathcal{B}^{(m)}_{\mathrm{score}}$ with $m>1$. What can we say otherwise? Some answers are given in the next section, where we argue that it is not (very) beneficial to only consider the single normal vector in $\mathrm{span}\{\Pi_{m}\}$.
4.3 Improving over Bhattacharyya/Pillai–Sinha-type bounds $\mathcal{B}^{(m)}_{\mathrm{score}}$
We start by noting the following, most of which we already met earlier.
Setup and notation.
Fix $m\geq 1$. Let $s_{\theta}\in L^{2}(\mu)$ be the square-root embedding of the model and write $\widetilde{Y}_{j}=Y_{j}\,s_{\theta},\quad\eta_{j}=\partial_{\theta}^{j}s_{\theta},\quad j\geq 1.$ Consider the jet subspace and the score-image span
$\mathcal{T}_{m}=\operatorname{span}\{\eta_{1},\dots,\eta_{m}\},\qquad\widetilde{\mathcal{T}}_{m}=\operatorname{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{m}\}.$
For the purposes of this section, introduce compact notation for the projection operators $^{\top}:=\operatorname{Proj}_{\mathcal{T}_{m}}, P:=P$$^{\perp}:=\operatorname{Proj}_{\mathcal{T}_{m}^{\perp}}$.
By Faà di Bruno/Bell polynomials we have the jet decomposition
$\eta_{j}\;=\;\tfrac{1}{2}\,\widetilde{Y}_{j}\;+\;W_{j},\qquad j\geq 2,$ (17)
where each remainder $W_{j}$ is a finite linear combination of products (monomials) in the lower-order $\widetilde{Y}_{\ell}$’s ($\ell<j$). Set
$W_{j}^{\perp}:=P^{\perp}W_{j},\qquad j\geq 2,$
the normal part of the remainder. Recall the normal vector
$\Pi_{m}:=P^{\perp}\eta_{m+1}\in\mathcal{T}_{m}^{\perp},$
and set the basic normal span
$\mathcal{N}_{m}=\operatorname{span}\{\Pi_{m}\}.$
Augmented normal span.
To capture normal directions produced by products in the Bell expansion define, for $m\geq 2$, the augmented normal span
$\mathcal{N}_{m}^{\mathrm{aug}}\;:=\;\operatorname{span}\{\Pi_{m},\ W_{2}^{\perp},\dots,W_{m}^{\perp}\}\;\subseteq\;\mathcal{T}_{m}^{\perp}.$ (18)
(If all $W_{j}^{\perp}$ are collinear with $\Pi_{m}$ then $\mathcal{N}_{m}^{\mathrm{aug}}=\mathcal{N}_{m}$.)
Lemma 6 (Containment of score span).
For every $m\geq 2$, we have the subspace inclusion
$\widetilde{\mathcal{T}}_{m}\;\subseteq\;\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}.$ (19)
Proof.
From (17) we get $\widetilde{Y}_{j}=2\eta_{j}-2W_{j}$. Apply the orthogonal decomposition relative to $\mathcal{T}_{m}$:
$\widetilde{Y}_{j}=P^{\top}\widetilde{Y}_{j}+P^{\perp}\widetilde{Y}_{j}=P^{\top}\widetilde{Y}_{j}-2W_{j}^{\perp},$
since $P^{\perp}\eta_{j}=0$ for $j\leq m$. It is immediate that $\widetilde{Y}_{j}\in\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}$ for every $j\leq m$ (of course, $\widetilde{Y}_{1}\in\mathcal{T}_{m},\,\forall m\geq 1$), and linear combinations of the $\widetilde{Y}_{j}$’s (i.e. $\widetilde{\mathcal{T}}_{m}$) belong to $\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}$, proving the containment. ∎
Proposition 7 (Universal comparison with augmented normal span).
Let $\widetilde{Z}_{0}=Z_{0}s_{\theta}\in L^{2}(\mu)$, with notation as before. Then, for $m\geq 2$,
$\mathcal{C}_{\mathrm{aug}}^{(m)}\;:=\;\big\|P^{\top}\widetilde{Z}_{0}\big\|^{2}\;+\;\big\|\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\big\|^{2}\;=\;\big\|\operatorname{Proj}_{\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\big\|^{2}\;\geq\;\big\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{m}}\widetilde{Z}_{0}\big\|^{2}\;=\;\mathcal{B}_{\mathrm{score}}^{(m)}.$ (20)
Equality holds iff $\operatorname{Proj}_{(\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}})\ominus\widetilde{\mathcal{T}}_{m}}\widetilde{Z}_{0}=0$.
Proof.
Lemma 6 gives $\widetilde{\mathcal{T}}_{m}\subseteq\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}$. For any closed subspaces $V_{1}\subseteq V_{2}$ and any vector $Z$ we have $\|\operatorname{Proj}_{V_{2}}Z\|\geq\|\operatorname{Proj}_{V_{1}}Z\|$. Taking $V_{1}=\widetilde{\mathcal{T}}_{m}$ and $V_{2}=\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}$ yields the inequality. Orthogonality of $\mathcal{T}_{m}$ and $\mathcal{N}_{m}^{\mathrm{aug}}$ gives the sum-of-squares decomposition on the left, and the equality criterion follows from standard projection geometry. ∎
Proposition 8 (Relation between basic and augmented spans; strict improvement).
With notation as above:
  1. 1.
    $\mathcal{N}_{m}^{\mathrm{aug}}=\mathcal{N}_{m}$ exactly when $W_{j}^{\perp}\in\mathcal{N}_{m}$ for every $j=2,\dots,m$. In particular, if all $W_{j}^{\perp}$ are collinear with $\Pi_{m}$, then the augmented span reduces to $\mathcal{N}_{m}$.
  2. 2.
    If $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{m}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{m}$, then strict improvement holds:
    $\mathcal{C}_{\mathrm{aug}}^{(m)}\;>\;\mathcal{B}_{\mathrm{score}}^{(m)}.$
  3. 3.
    Under the condition in point 1 (so $\mathcal{N}_{m}^{\mathrm{aug}}=\mathcal{N}_{m}$), the comparison (20) reduces to
    $\mathcal{C}^{(m)}\;=\;\big\|P^{\top}\widetilde{Z}_{0}\big\|^{2}+\big\|\operatorname{Proj}_{\mathcal{N}_{m}}\widetilde{Z}_{0}\big\|^{2}\;\geq\;\mathcal{B}_{\mathrm{score}}^{(m)},$
    with strict inequality whenever $\operatorname{Proj}_{\mathcal{N}_{m}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{m}\not\subseteq\widetilde{\mathcal{T}}_{m}$.
Proof.
(i) is immediate from the definition $\mathcal{N}_{m}^{\mathrm{aug}}=\operatorname{span}\{\Pi_{m},W_{2}^{\perp},\dots,W_{m}^{\perp}\}$.
(ii) By Proposition 7, $\mathcal{C}_{\mathrm{aug}}^{(m)}=\|\operatorname{Proj}_{\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\|^{2}\geq\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{m}}\widetilde{Z}_{0}\|^{2}$. If $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{m}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{m}$, then the orthogonal complement $(\mathcal{T}_{m}\oplus\mathcal{N}_{m}^{\mathrm{aug}})\ominus\widetilde{\mathcal{T}}_{m}$ is nontrivial and contains a component of $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}$; therefore projection onto the larger space strictly increases the squared norm and $\mathcal{C}_{\mathrm{aug}}^{(m)}>\mathcal{B}_{\mathrm{score}}^{(m)}$.
(iii) follows by substituting $\mathcal{N}_{m}^{\mathrm{aug}}=\mathcal{N}_{m}$ into Proposition 7 and repeating the strictness argument as in (ii). ∎
Remark 4.
It is trivial to see, from the proof of Lemma 6, that we get $\mathcal{C}_{\mathrm{aug}}^{(m)}\;\geq\;\mathcal{B}_{\mathrm{score}}^{(m)}$ (and strict inequality with sufficient conditions above) with the augmented normal space being defined as $\operatorname{span}\{W_{2}^{\perp},\dots,W_{m}^{\perp}\}\;\subseteq\;\mathcal{T}_{m}^{\perp},$ without the $\Pi_{m}$ (cf. (18)). Thus we are not “cheating" by considering a higher-order derivative in the comparison.
4.4 Summary and comparison with Bhattacharyya-type bounds
The classical Bhattacharyya inequality ([8]; see also [9, 10]) generalizes the CRB by incorporating higher-order derivatives of the model density. The core idea here, as presented by Pillai and Sinha (cf. [11]), can be systematically understood within a projection framework in $L^{2}(P_{\theta})$. In their construction, one considers the span $\widetilde{\mathcal{T}}_{m}=\operatorname{span}\{\widetilde{Y}_{1},\dots,\widetilde{Y}_{m}\}$ of score derivatives (as mentioned previously, they do not exactly consider the score derivatives but the concept remains the same) and projects the centered estimator residual $\widetilde{Z}_{0}=(T-\theta)s_{\theta}$ onto $\widetilde{\mathcal{T}}_{m}$. This leads to an explicit sequence of bounds
$\mathcal{B}_{\mathrm{score}}^{(m)}\;=\;\|\mathrm{Proj}_{\widetilde{\mathcal{T}}_{m}}\widetilde{Z}_{0}\|^{2},$
which coincides with bounds of the kind obtained by Bhattacharyya, but now using derivatives of the score (not $\frac{\partial^{k}_{\theta}f}{f}$), and recovers the CRB at $m=1$.
In contrast, we obtained further refinements by embedding the statistical model manifold extrinsically into the Hilbert space $L^{2}(\mu)$ and exploiting the associated differential geometry. For $m=1$, we showed that the CRB is improved by an explicit curvature correction term involving the second fundamental form $\Pi(\eta_{1},\eta_{1})$, where $\eta_{1}$ is the score direction (see Theorem 2). This yields a strictly sharper bound whenever the estimator residual has nontrivial projection in the normal direction of the model manifold – considerations of this kind are not made in the Bhattacharyya-type bounds.
For $m>1$, we constructed the jet space $\mathcal{T}_{m}=\mathrm{span}\{\eta_{1},\dots,\eta_{m}\}$ of embedded derivatives of the statistical model, and defined the augmented normal span
$\mathcal{N}^{\mathrm{aug}}_{m}=\mathrm{span}\big\{\Pi_{m},\,W_{2}^{\perp},\dots,W_{m}^{\perp}\big\},$
where $\Pi_{m}$ encodes curvature (“beyond order mm") and the $W_{j}^{\perp}$ are additional orthogonal components arising from Faà di Bruno expansions of the embedding. Our refined bound then reads
$\mathcal{C}_{\mathrm{aug}}^{(m)}\;=\;\big\|\mathrm{Proj}_{\mathcal{T}_{m}\oplus\mathcal{N}^{\mathrm{aug}}_{m}}\,\widetilde{Z}_{0}\big\|^{2},$
and we established that
$\mathcal{C}_{\mathrm{aug}}^{(m)}\;\geq\;\mathcal{B}_{\mathrm{score}}^{(m)},$
with strict inequality in generic curved models (see sufficient conditions stated in Proposition 8). Thus, the classical Bhattacharyya-type inequality/Pillai–Sinha sequence may be viewed as a score-only case, while our extrinsic geometric refinement yields systematically sharper bounds by incorporating curvature and normal components that lie beyond the score span (and one may improve upon these classical bounds even when only the curvature part is ignored; see Remark 4). These also capture the correct geometric picture of the setting under consideration. What we have shown may be succintly stated this way: Defining, for $m\geq 1$, $\mathcal{B}_{\mathrm{jet}}^{(m)}\;=\;\|\mathrm{Proj}_{\mathcal{T}_{m}}\widetilde{Z}_{0}\|^{2},$ we get the sequence of bounds $\mathcal{B}_{\mathrm{jet}}^{(1)}\leq\mathcal{B}_{\mathrm{jet}}^{(1)}\leq\mathcal{B}_{\mathrm{jet}}^{(3)}\leq\dots$ ($\mathcal{B}_{\mathrm{jet}}^{(1)}$ equalling the CRB), with
$\mathcal{B}_{\mathrm{jet}}^{(m+1)}-\mathcal{B}_{\mathrm{jet}}^{(m)}=\mathcal{C}^{(m)},$
where $\mathcal{C}^{(m)}$ is precisely the curvature correction beyond order $m$; see (16).
To conclude, our extrinsic geometry point of view results in strict improvements over classical bounds whenever the estimator error does not purely lie in the span of the score and its derivatives. This is not the perspective taken by Bhattacharyya [8]. One may argue that second-order information, as present in the curvature via second fundamental form, is taken into account in the second-order Bhattacharyya-type bounds, but this does not correspond to the curvature in any proper mathematical sense.
Some of the ideas developed in this section will be illustrated in the analytical Example 3 below.
5 Analytical examples
We illustrate Theorems 2, 5, and some of the key results stated in Section 4.3 with two fully analytical examples below; both explicitly demonstrate improvement over the CRB (i.e. $\mathcal{B}^{(1)}_{\text{score}}=\mathcal{B}^{(1)}_{\text{jet}}\equiv\mathcal{B}^{(1)}$), while the second also illuminates results comparing $\mathcal{B}_{\mathrm{score}}^{(m)}$ with our jet subspace-based curvature correction theorems.
Example 2 (One-parameter curved normal family).
We consider the family of normal laws, where $P_{\theta}$ is the distribution
$\mathcal{N}\!\bigl(\mu(\theta),\,\sigma^{2}(\theta)\bigr),\qquad\mu(\theta)=\theta,\qquad\sigma^{2}(\theta)=1+\theta^{2},\qquad\theta\in\mathbb{R},$
with Lebesgue base measure $\mu$ and strictly positive density
$f(x;\theta)=\frac{1}{\sqrt{2\pi(1+\theta^{2})}}\,\exp\!\left(-\frac{(x-\theta)^{2}}{2(1+\theta^{2})}\right).$
Write $B:=1+\theta^{2}$ and $U:=X-\theta$ so that, under $P_{\theta}$, we have $U\sim\mathcal{N}(0,B)$. The score and its first derivative are
$Y_{1}(x;\theta)$ $=\partial_{\theta}\log f(x;\theta)=-\frac{\theta}{B}+\frac{U}{B}+\frac{\theta U^{2}}{B^{2}},$ (21)
$Y_{2}(x;\theta)$ $=\partial_{\theta}^{2}\log f(x;\theta)=-\frac{2}{B^{2}}-\frac{4\theta}{B^{2}}U+\frac{1-3\theta^{2}}{B^{3}}U^{2}.$ (22)
Recall the jet fields
$\eta_{1}=\partial_{\theta}s_{\theta}=\frac{1}{2}\,Y_{1}\,s_{\theta},\qquad\eta_{2}=\partial_{\theta}^{2}s_{\theta}=\frac{1}{2}\Bigl(Y_{2}+\tfrac{1}{2}Y_{1}^{2}\Bigr)s_{\theta}.$
Consider an unbiased estimator $T(X)$ of $\theta$. Set $Z_{0}=T(X)-\theta$ and identify $Z_{0}s_{\theta}\in L^{2}(\mu)$. For the location-equivariant unbiased estimator $T(X)=X$, we have $Z_{0}=U$ and $\operatorname{Var}_{\theta}[T]=\mathbb{E}_{\theta}[U^{2}]=B$.
Using $U\sim\mathcal{N}(0,B)$ and the Gaussian moment formulae $\mathbb{E}_{\theta}[U]=0$, $\mathbb{E}_{\theta}[U^{2}]=B$, $\mathbb{E}_{\theta}[U^{4}]=3B^{2}$, $\mathbb{E}_{\theta}[U^{6}]=15B^{3}$, $\mathbb{E}_{\theta}[U^{8}]=105B^{4}$, one obtains:
$\mathcal{I}(\theta)=\mathbb{E}_{\theta}[Y_{1}^{2}]$ $=\frac{1+3\theta^{2}}{B^{2}},$ (23)
$\mathbb{E}_{\theta}[Y_{1}Y_{2}]$ $=-\,\frac{2\theta(1+5\theta^{2})}{B^{3}},$ (24)
$\mathbb{E}_{\theta}[Y_{2}^{2}]$ $=\frac{3+10\theta^{2}+43\theta^{4}}{B^{4}},$ (25)
$\mathbb{E}_{\theta}[Y_{1}^{3}]$ $=\frac{6\theta+14\theta^{3}}{B^{3}},$ (26)
$\mathbb{E}_{\theta}[U\,Y_{2}]$ $=-\,\frac{4\theta}{B},$ (27)
$\mathbb{E}_{\theta}[U\,Y_{1}^{2}]$ $=\frac{4\theta}{B}.$ (28)
The Gram matrix $G\in\mathbb{R}^{2\times 2}$ has entries $G_{ij}=\langle\eta_{i},\eta_{j}\rangle$:
$G_{11}$ $=\left\langle\frac{1}{2}Y_{1}s_{\theta},\frac{1}{2}Y_{1}s_{\theta}\right\rangle=\frac{1}{4}\,\mathbb{E}_{\theta}[Y_{1}^{2}]=\frac{1+3\theta^{2}}{4B^{2}},$ (29)
$G_{12}$ $=\left\langle\eta_{1},\eta_{2}\right\rangle=\frac{1}{4}\left(\mathbb{E}_{\theta}[Y_{1}Y_{2}]+\frac{1}{2}\mathbb{E}_{\theta}[Y_{1}^{3}]\right)=\frac{\theta(1-3\theta^{2})}{4B^{3}},$ (30)
$G_{22}$ $=\left\langle\eta_{2},\eta_{2}\right\rangle=\frac{1}{4}\left(\mathbb{E}_{\theta}[Y_{2}^{2}]+\mathbb{E}_{\theta}[Y_{2}Y_{1}^{2}]+\frac{1}{4}\mathbb{E}_{\theta}[Y_{1}^{4}]\right)=\frac{19+34\theta^{2}+75\theta^{4}}{16\,B^{4}},$ (31)
where we used, in addition to (25), the identities
$\mathbb{E}_{\theta}[Y_{2}Y_{1}^{2}]=\frac{-55\theta^{4}-18\theta^{2}+1}{B^{4}},\qquad\mathbb{E}_{\theta}[Y_{1}^{4}]=\frac{3\,(41\theta^{4}+22\theta^{2}+1)}{B^{4}}.$
For $Z_{0}s_{\theta}$, the coefficients against $\eta_{1},\eta_{2}$ are
$a_{1}$ $:=\langle Z_{0}s_{\theta},\eta_{1}\rangle=\frac{1}{2}\,\mathbb{E}_{\theta}[Z_{0}Y_{1}]=\frac{1}{2},$ (32)
$a_{2}$ $:=\langle Z_{0}s_{\theta},\eta_{2}\rangle=\frac{1}{2}\left(\mathbb{E}_{\theta}[Z_{0}Y_{2}]+\frac{1}{2}\mathbb{E}_{\theta}[Z_{0}Y_{1}^{2}]\right)=-\frac{\theta}{B},$ (33)
where we used $\mathbb{E}_{\theta}[Z_{0}Y_{1}]=1$, and (27)–(28) with $Z_{0}=U$.
The CRB ($m=1$) reads
$\mathcal{B}^{(1)}_{\text{score}}=\mathcal{B}^{(1)}_{\text{jet}}\equiv\mathcal{B}^{(1)}=\frac{1}{\mathcal{I}(\theta)}=\frac{B^{2}}{1+3\theta^{2}}=\frac{(1+\theta^{2})^{2}}{1+3\theta^{2}}.$
The Christoffel coefficient is
$\Gamma_{11}=\frac{\langle\eta_{2},\eta_{1}\rangle}{\langle\eta_{1},\eta_{1}\rangle}=\frac{G_{12}}{G_{11}}=\frac{-3\theta^{3}+\theta}{3\theta^{4}+4\theta^{2}+1}.$
The second fundamental form is the normal component
$\mathrm{II}(\eta_{1},\eta_{1})=\eta_{2}-\Gamma_{11}\,\eta_{1}\;\in\;\mathcal{T}_{1}^{\perp}.$
Its squared norm and correlation with $Z_{0}s_{\theta}$ are
$\|\mathrm{II}(\eta_{1},\eta_{1})\|^{2}$ $=G_{22}-\Gamma_{11}^{2}G_{11}=\frac{189\theta^{6}+201\theta^{4}+87\theta^{2}+19}{48\theta^{10}+208\theta^{8}+352\theta^{6}+288\theta^{4}+112\theta^{2}+16},$ (34)
$\langle Z_{0}s_{\theta},\mathrm{II}(\eta_{1},\eta_{1})\rangle$ $=a_{2}-\Gamma_{11}a_{1}=-\frac{3\theta}{2\,(3\theta^{2}+1)}.$ (35)
Therefore, our curvature-corrected lower bound for $m=1$ is
$\operatorname{Var}_{\theta}[T]\;\geq\;\mathcal{B}^{(1)}+\frac{\langle Z_{0}s_{\theta},\mathrm{II}(\eta_{1},\eta_{1})\rangle^{2}}{\|\mathrm{II}(\eta_{1},\eta_{1})\|^{2}}=\frac{B^{2}}{1+3\theta^{2}}+\frac{36\,\theta^{2}\left(\theta^{8}+4\theta^{6}+6\theta^{4}+4\theta^{2}+1\right)}{567\theta^{8}+792\theta^{6}+462\theta^{4}+144\theta^{2}+19},$ (36)
and we get a strictly positive curvature correction term for all $\theta\neq 0$. Of course, when $\theta=0$, $\operatorname{Var}_{\theta}[T]=\mathcal{B}^{(1)}=1$ already.
Example 3 (Strict gains at $m=1$ and $m=2$).
Let $s_{0}(x)=(2\pi)^{-1/4}e^{-x^{2}/4}$ be the Gaussian square root and let $H_{n}$ be the probabilists’ Hermite polynomials with $\langle H_{m}s_{0},H_{n}s_{0}\rangle=\mathbb{E}_{\varphi}[H_{m}H_{n}]=n!\,\delta_{mn}$ under $\varphi\sim\mathcal{N}(0,1)$.
Consider the one-parameter square-root family
$s_{\theta}(x)\;=\;s_{0}(x)\,\exp\!\Big(a_{1}(\theta)\,H_{1}(x)+a_{2}(\theta)\,H_{2}(x)\Big),\qquad\theta\in\mathbb{R},$ (37)
with
$a_{1}(\theta)=\theta,\qquad a_{2}(\theta)=\theta+\tfrac{c}{2}\,\theta^{2},$
where $c\in\mathbb{R}\setminus\{0\}$ is a fixed constant. Write $u(\theta):=\tfrac{1}{2}\log f(\cdot;\theta)=a_{1}(\theta)H_{1}+a_{2}(\theta)H_{2}+\log s_{0}$ so that
$Y_{1}=2\,\partial_{\theta}u,\qquad Y_{2}=2\,\partial_{\theta}^{2}u,\qquad\widetilde{Y}_{k}:=Y_{k}s_{\theta},\quad k=1,2.$
At the base point $\theta=0$, we have
$a_{1}^{\prime}(0)=1,\quad a_{2}^{\prime}(0)=1,\quad a_{2}^{\prime\prime}(0)=c,\qquad Y_{1}(0)=2(H_{1}+H_{2}),\quad Y_{2}(0)=2c\,H_{2}.$
The jets are
$\eta_{1}=\partial_{\theta}s_{\theta}=\tfrac{1}{2}Y_{1}s_{\theta},\qquad\eta_{2}=\partial_{\theta}^{2}s_{\theta}=\tfrac{1}{2}\Big(Y_{2}+\tfrac{1}{2}Y_{1}^{2}\Big)s_{\theta}.$
Evaluating at $\theta=0$ and expanding products with the Hermite product identity
$H_{m}H_{n}=\sum_{k=0}^{\min(m,n)}k!\binom{m}{k}\binom{n}{k}H_{m+n-2k}$
yields
$\eta_{1}(0)$ $=(H_{1}+H_{2})\,s_{0},$
$Y_{1}(0)^{2}$ $=4(H_{1}+H_{2})^{2}=4\big(H_{1}^{2}+2H_{1}H_{2}+H_{2}^{2}\big)$
$=4\big[(H_{3}+2H_{1})+2(H_{3}+2H_{1})+(H_{4}+4H_{2}+2)\big]$
$=12H_{3}+24H_{1}+16H_{2}+4H_{4}+8,$
$\eta_{2}(0)$ $=\tfrac{1}{2}\Big(2c\,H_{2}+\tfrac{1}{2}Y_{1}(0)^{2}\Big)s_{0}=s_{0}\Big(H_{4}+2H_{3}+(5+c)H_{2}+4H_{1}+3H_{0}\Big).$
The score-image span at order $m=2$ is
$\widetilde{\mathcal{T}}_{2}(0)=\operatorname{span}\{\widetilde{Y}_{1}(0),\widetilde{Y}_{2}(0)\}=\operatorname{span}\{2(H_{1}+H_{2})s_{0},\;2c\,H_{2}s_{0}\}=\operatorname{span}\{H_{1}s_{0},\;H_{2}s_{0}\}.$
The $2$-jet span is
$\mathcal{T}_{2}(0)=\operatorname{span}\{\eta_{1}(0),\eta_{2}(0)\}=\operatorname{span}\Big\{(H_{1}+H_{2})s_{0},\;\big(H_{4}+2H_{3}+(5+c)H_{2}+4H_{1}+3H_{0}\big)s_{0}\Big\}.$
Thus $\mathcal{T}_{2}(0)$ is a two-dimensional subspace whose projection onto $\widetilde{\mathcal{T}}_{2}(0)$ is
$\operatorname{Proj}_{\widetilde{\mathcal{T}}_{2}(0)}\eta_{1}(0)=(H_{1}+H_{2})s_{0},\qquad\operatorname{Proj}_{\widetilde{\mathcal{T}}_{2}(0)}\eta_{2}(0)=\big(4H_{1}+(5+c)H_{2}\big)s_{0},$
which are linearly independent for all $c\neq-1$. The normal component of $\eta_{2}(0)$ relative to $\widetilde{\mathcal{T}}_{2}(0)$ is
$\operatorname{Proj}_{\widetilde{\mathcal{T}}_{2}(0)^{\perp}}\eta_{2}(0)\;=\;s_{0}\big(2H_{3}+H_{4}+3H_{0}\big),$
which is nonzero; thus $\eta_{2}(0)$ causes $\mathcal{T}_{2}(0)$ to be tilted in the directions in $\widetilde{\mathcal{T}}_{2}(0)^{\perp}$ and is precisely the geometric source of the curvature correction below.
Define the unbiased estimator $T(X)=H_{3}(X)$. Then $Z_{0}:=T-\mathbb{E}_{\theta}[H_{3}(X)]$ satisfies, at $\theta=0$,
$Z_{0}\,s_{0}\;=\;H_{3}s_{0}.$
At $m=1$, the score span is $\widetilde{\mathcal{T}}_{1}(0)=\operatorname{span}\{\widetilde{Y}_{1}(0)\}=\operatorname{span}\{(H_{1}{+}H_{2})s_{0}\}$, hence
$\mathcal{B}^{(1)}_{\text{score}}=\mathcal{B}^{(1)}_{\text{jet}}\equiv\mathcal{B}^{(1)}=\big\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{1}(0)}(H_{3}s_{0})\big\|^{2}=0,$
since $H_{3}\perp H_{1},H_{2}$. Recall again that
$\eta_{1}=(H_{1}+H_{2})s_{0},\qquad\eta_{2}=(H_{4}+2H_{3}+(5+c)H_{2}+4H_{1}+3H_{0})s_{0}.$
The projection coefficient is
$\Gamma_{11}\;=\;\frac{\langle\eta_{2},\eta_{1}\rangle}{\langle\eta_{1},\eta_{1}\rangle}=\frac{14+2c}{3},$
so that
$\Pi_{1}\;=\;\eta_{2}-\Gamma_{11}\eta_{1}=s_{0}\Big(H_{4}+2H_{3}+\tfrac{1+c}{3}H_{2}-\tfrac{2(1+c)}{3}H_{1}+3H_{0}\Big).$
The inner product with $H_{3}s_{0}$ is
$\langle H_{3}s_{0},\Pi_{1}\rangle=2\cdot 3!=12,$
and the squared norm of $\Pi_{1}$ is
$\|\Pi_{1}\|^{2}=4!\cdot 1^{2}+3!\cdot 2^{2}+2!\left(\tfrac{1+c}{3}\right)^{2}+1!\left(\tfrac{2(1+c)}{3}\right)^{2}+0!\cdot 3^{2}=57+\tfrac{2}{3}(1+c)^{2}.$
Therefore the curvature contribution at order $1$ is
$\|\mathrm{Proj}_{\operatorname{span}\{\Pi_{1}\}}(H_{3}s_{0})\|^{2}=\frac{\langle H_{3}s_{0},\Pi_{1}\rangle^{2}}{\|\Pi_{1}\|^{2}}=\frac{144}{\,57+\tfrac{2}{3}(1+c)^{2}\,}.$
In particular, we get a strictly positive improvement
$\mathcal{C}^{(1)}\;=\;\mathcal{B}^{(1)}\;+\;\frac{144}{\,57+\tfrac{2}{3}(1+c)^{2}\,}$,
which is strictly less than the full variance $\|H_{3}s_{0}\|^{2}=6$ (since $\Pi_{1}$ is not parallel to $H_{3}s_{0}$ for any real $c$). For example, at $c=1$ this evaluates to $\approx 2.41$.
At $m=2$, $\widetilde{\mathcal{T}}_{2}(0)=\operatorname{span}\{H_{1}s_{0},H_{2}s_{0}\}$ as above, so still
$\mathcal{B}^{(2)}_{\text{score}}=\big\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{2}(0)}(H_{3}s_{0})\big\|^{2}=0.$
It is easy to see that $\mathcal{B}^{(2)}_{\mathrm{jet}}=\mathcal{C}^{(1)}$ always; that is, this is not example dependent. Therefore, also in our example,
$\mathcal{B}^{(2)}_{\mathrm{jet}}\;=\;\|\operatorname{Proj}_{\mathcal{T}_{2}}(H_{3}s_{0})\|^{2}\;=\;\frac{432}{\,2c^{2}+4c+173\,}.$
In particular, $\mathcal{B}^{(2)}_{\mathrm{jet}}>0$ for every real $c$. Since $\mathcal{B}^{(2)}_{\text{score}}=0$ and $\mathcal{C}^{(2)}\geq\mathcal{B}^{(2)}_{\mathrm{jet}}>0$ for all $c\in\mathbb{R}$ (of course, we are concerned with $c\neq 0$ in this example), we have demonstrated strict improvement also at $m=2$ with Theorem 5.
Explicitly, we have
$\eta_{3}/s_{0}=(H_{1}+H_{2})^{3}+3c\,(H_{1}H_{2}+H_{2}^{2})=H_{6}+3H_{5}+(15+3c)H_{4}+(25+3c)H_{3}$
$(45+12c)H_{2}+(33+6c)H_{1}+(14+6c)H_{0}.$
The tangent space at order $m=2$ is $\mathcal{T}_{2}=\operatorname{span}\{\eta_{1},\eta_{2}\}$, and the second fundamental vector is
$\mathrm{II}_{2}\;=\;\eta_{3}-\operatorname{Proj}_{\mathcal{T}_{2}}\eta_{3}.$
The residual variance bound at order $m=2$ is then given by
$\mathcal{C}^{(2)}=\big\|\operatorname{Proj}_{\mathcal{T}_{2}}(H_{3}s_{0})\big\|^{2}+\big\|\operatorname{Proj}_{\operatorname{span}\{\mathrm{II}_{2}\}}(H_{3}s_{0})\big\|^{2}.$
Recall
$\mathcal{C}^{(1)}=\frac{432}{\,2c^{2}+4c+173\,}$,
and a direct symbolic computation yields
$\mathcal{C}^{(2)}-\mathcal{C}^{(1)}\;=\;\frac{18\bigl(6c^{3}+38c^{2}-209c+65\bigr)^{2}}{(2c^{2}+4c+173)\,\bigl(306c^{4}+2328c^{3}+9349c^{2}+16070c+216283\bigr)}.$
Substituting $c=1$, we obtain
$\mathcal{C}^{(2)}-\mathcal{C}^{(1)}\approx 0.0041.$
The improvement is minimal for this particular value of $c$. Can we better this significantly using the stronger Proposition 8?
The sufficient conditions stated in Proposition 8 can be checked in this example with $m=2$: If $\operatorname{Proj}_{\mathcal{N}_{m}^{\mathrm{aug}}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{m}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{m}$, then strict improvement holds:
$\mathcal{C}_{\mathrm{aug}}^{(m)}\;>\;\mathcal{B}_{\mathrm{score}}^{(m)}.$
Towards this, recall the Bell remainder for $\eta_{2}$:
$\eta_{2}\;=\;\tfrac{1}{2}\widetilde{Y}_{2}+W_{2},\qquad W_{2}\;=\;\tfrac{1}{4}Y_{1}^{2}s_{0}=s_{0}\big(2+6H_{1}+4H_{2}+3H_{3}+H_{4}\big).$
Decompose $W_{2}$ into its projection onto $\mathcal{T}_{2}$ and the normal (residual) part
$W_{2}=\operatorname{Proj}_{\mathcal{T}_{2}}W_{2}\;+\;W_{2}^{\perp},\qquad W_{2}^{\perp}:=\operatorname{Proj}_{\mathcal{T}_{2}^{\perp}}W_{2}.$
We want to find the residual
$W_{2}^{\perp}\;=\;W_{2}-\operatorname{Proj}_{\mathcal{T}_{2}}(W_{2}),\qquad\mathcal{T}_{2}=\operatorname{span}\{\eta_{1},\eta_{2}\}.$
Using $\langle H_{m}s_{0},H_{n}s_{0}\rangle=\delta_{mn}\,n!$, we get
$\|W_{2}\|^{2}$ $=2^{2}\cdot 0!+6^{2}\cdot 1!+4^{2}\cdot 2!+3^{2}\cdot 3!+1^{2}\cdot 4!$
$=4+36+32+54+24=150,$
$\|\eta_{1}\|^{2}$ $=1^{2}\cdot 1!+1^{2}\cdot 2!=1+2=3,$
$\|\eta_{2}\|^{2}$ $=3^{2}\cdot 0!+4^{2}\cdot 1!+(5+c)^{2}\cdot 2!+2^{2}\cdot 3!+1^{2}\cdot 4!$
$=9+16+2(5+c)^{2}+24+24=2c^{2}+20c+123,$
$\langle\eta_{1},\eta_{2}\rangle$ $=4+2(5+c)=14+2c,$
$\langle W_{2},\eta_{1}\rangle$ $=6\cdot 1!+4\cdot 2!=6+8=14,$
$\langle W_{2},\eta_{2}\rangle$ $=2\cdot 3\cdot 0!+6\cdot 4\cdot 1!+4(5+c)\cdot 2!+3\cdot 2\cdot 3!+1\cdot 1\cdot 4!$
$=6+24+8(5+c)+36+24=130+8c.$
Write $\operatorname{Proj}_{\mathcal{T}_{2}}(W_{2})=a\,\eta_{1}+b\,\eta_{2}$. The coefficients $(a,b)$ solve
$\begin{pmatrix}\|\eta_{1}\|^{2}&\langle\eta_{1},\eta_{2}\rangle\\[2.15277pt] \langle\eta_{1},\eta_{2}\rangle&\|\eta_{2}\|^{2}\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}=\begin{pmatrix}\langle W_{2},\eta_{1}\rangle\\ \langle W_{2},\eta_{2}\rangle\end{pmatrix}$,
i.e.
$\begin{pmatrix}3&14+2c\\[4.0pt] 14+2c&2c^{2}+20c+123\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}=\begin{pmatrix}14\\ 130+8c\end{pmatrix}$.
One obtains, after simplification, the closed form
$W_{2}^{\perp}\;=\;s_{0}\Bigg[\frac{4\big(c^{2}+5c-59\big)}{\,2c^{2}+4c+173\,}\,H_{0}\;+\;\frac{12\big(11c+30\big)}{\,2c^{2}+4c+173\,}\,H_{1}\;-\;\frac{6\big(11c+30\big)}{\,2c^{2}+4c+173\,}\,H_{2}\;$
$\frac{6c^{2}+20c+131}{\,2c^{2}+4c+173\,}\,H_{3}\;+\;\frac{2c^{2}+8c-21}{\,2c^{2}+4c+173\,}\,H_{4}\Bigg],$
and
$\|W_{2}^{\perp}\|^{2}\;=\;\frac{2\big(82c^{2}+428c+1051\big)}{\,2c^{2}+4c+173\,}>0.$
We note that the coefficient of $H_{3}$ in $W_{2}^{\perp}$, which is $\frac{6c^{2}+20c+131}{\,2c^{2}+4c+173\,}$, is not zero for any $c\in\mathbb{R}$. With $\mathcal{N}_{2}^{\mathrm{aug}}=\mathrm{span}\{\Pi_{2},W_{2}^{\perp}\}$ and $\widetilde{\mathcal{T}}_{2}=\mathrm{span}\{H_{1}s_{0},H_{2}s_{0}\}$, this implies that the sufficient conditions $\operatorname{Proj}_{\mathcal{N}_{2}^{\mathrm{aug}}}H_{3}s_{0}\neq 0$ and $\mathcal{N}_{2}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{2}$ are satisfied, and Proposition 8 guarantees that
$\mathcal{C}_{\mathrm{aug}}^{(2)}\;>\;\mathcal{B}_{\mathrm{score}}^{(2)}.$
Using explicit computations, we have
$\mathcal{C}^{(2)}_{\mathrm{aug}}\;\geq\;\big\|\operatorname{Proj}_{\mathcal{T}_{2}}(H_{3}s_{0})\big\|^{2}\;+\;\frac{\langle H_{3}s_{0},W_{2}^{\perp}\rangle^{2}}{\|W_{2}^{\perp}\|^{2}}\;=\;\frac{18\,(18c^{2}+84c+245)}{\,82c^{2}+428c+1051\,}$,
which approximately evaluates to $4.001$ at $c=1$, significantly improving over $\mathcal{C}^{(1)}$ and $\mathcal{C}^{(2)}$ (both of which were approximately $2.41$); in fact, one can establish that $\mathcal{C}^{(2)}_{\mathrm{aug}}>\mathcal{C}^{(1)}$ for all $c$. Note that the lower bound in the last display is due to the fact that we have not included $\Pi_{2}$ in computing the second term and, in fact, $\mathcal{N}_{2}^{\mathrm{aug}}=\mathrm{span}\{\Pi_{2},W_{2}^{\perp}\}$.
6 Conclusion
In this work we developed refinements of the classical Cramér-Rao lower bound, and some of its own refinements such as the Bhattacharyya-type bounds that use higher-order score functions, in a manner consistent with the local differential geometry of the statistical model. The CRB case incorporated curvature-aware corrections derived from the second fundamental form on tangent vectors of the statistical manifold embedded into a fixed Hilbert space using the square root embedding, while the higher-order version also considered the decomposition of the corresponding jets in terms of the true scores using a special case of the Faà di Bruno formula. We detailed explicit examples where the geometric correction terms result in definite improvements over the CRB and its variants. This confirms that curvature-aware corrections, when properly projected and interpreted within a Riemannian framework, yield provable and practically relevant enhancements. Further research may examine the generality and limitations of these geometric corrections in broader model classes, such as those with higher-order nonlinearity in the statistical embedding or with weaker assumptions on regularity than what we allowed, for instance. The connection between extrinsic curvature and information-theoretic optimality, particularly beyond the second-order regime, remains an area with substantial theoretical and methodological potential.
Acknowledgements.
We are thankful to Prof. Pillai for sharing his paper [11], which formed the motivation for this work.